OpenStudy (anonymous):

2. Every day you collect 4 pennies and keep them in a jar. So the first day you have 4 pennies, the second day you have 8 pennies, the third day you have 12 pennies, etc. How many do you have on the 15th day? (1 point) 75 19 15 60

jimthompson5910 (jim_thompson5910):

plug in n = 15 into a(n) = 4 + 4(n-1) to get your answer

OpenStudy (anonymous):

4+ 4(14) =15(4)= 60

OpenStudy (anonymous):

4. To which set of numbers does the number belong? (1 point) square root 51 rational numbers irrational numbers integers natural numbers

OpenStudy (anonymous):

can someone help me

jimthompson5910 (jim_thompson5910):

is 51 a perfect square?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

it is? so there's some integer that you can square to get 51?

jimthompson5910 (jim_thompson5910):

ex: 100 is a perfect square since 10^2 = 100

jimthompson5910 (jim_thompson5910):

4 is a perfect square because 2^2 = 2*2 = 4